Aaliya:

Michael works in the marketing department of a company. He recorded the sales of his company for 10 consecutive months. The amount of sales, n(t), over time t, in months, can be modeled by a cubic function. Each of the following functions is a different form of the cubic model for the situation given above. Which form would be the most helpful if attempting to determine the number of months after which the sales were approximately constant?

mhanifa:

It is incomplete. Could you attach the full question or the missing part?

Vocaloid:

A) n(t) = 0.1(t - 8)3 + 51.2 B) n(t) = 0.1t2(t - 24) + 19.2t C) n(t) = 0.1t3 - 2.4t2 + 19.2t D) n(t) = 0.1(t3 - 24t2 + 192t)

Vocaloid:

Since it’s asking for the times at which sales are constant, we’re looking for the part of the graph where the derivative is 0 Think about which function would best let you determine that